So, to get you using L^{A}TEX :

Exercise 1. Find and List all the Keywords (Schlüsselwörter) for important concepts in operating systems that occur in Chapters 1-9 of Galvin/Silberschatz or Tannenbaum or equivalent, and arrange them in a hierarchy.

Classify them into • Data Structures (e.g. Files)

• Objects (e.g. Processes)

• Methods (e.g. Remote-Procedure Call)

• Features (e.g. Concurrency)

• Any other class you can justify

(with your justification) Organise them into a tree of subclasses (e.g. is RPC a form of communication?) Write your answer in L^{A}TEX Hint for Exercise 1: You can use itemize environments inside other itemize environments, like:

\begin{itemize}
\item One thing
\begin{itemize}
\item Lots of little things
\item More little things
\end{itemize}
\item One more big thing
\end{itemize}

This gives

• One thing

– Lots of little things

– More little things

• One more big thing Here’s how a document looks in L^{A}TEX

\documentclass{article}
\usepackage{tla,rawfonts}
\author{Not Me}
\title{A Short History of the World}
\begin{document}
\maketitle
\section{The Parable}
Once upon a time, there were no computers.
Life was
\begin{itemize}
\item Simple
\item Brutish
\item Short
\end{itemize}
rather like the people
\end{document}

Here’s (roughly) how it looks compiled:

A Short History of the World

Not Me

1 The Parable Once upon a time, there were no computers. Life was

• Simple

• Brutish

• Short rather like the people